# Symmetric Entropy-Constrained Video Coding for Machines

### v1

@@ III-B2 Bi-directional Entropy-Constraint Mechanism
Figure 5: The t-SNE visualization of intermediate features $f_{\textit{MVS}}^{1/8}$ in semantic decoder with and without symmetric information alignment. After alignment via the BiEC mechanism, intra-class distances are effectively reduced, and the intermediate features $f_{\textit{MVS}}^{1/8}$ in the decoding process exhibit more discriminative semantic representations.
To enforce symmetry, we further impose a conditional entropy constraint in the reverse direction, making $f_{\textit{VB}}$ a subset of $f_{\textit{MVS}}$. Through this bi-directional constraint, the features $f_{\textit{MVS}}$ in the semantic decoder and the visual-backbone-extracted representations $f_{\textit{VB}}$ are softly aligned, as illustrated in Fig. 4. The above process is formulated as follows:
@@
Specifically, each distribution estimation module contains a convolution layer and a ResBlock [21]. Estimated distribution parameters $\mu$ and $\sigma$ have the same shape as the corresponding feature $f$. Following existing methods [50, 29], the Laplace distribution is used to model the conditional probability of the features $f_{\textit{MVS}}$ and $f_{\textit{VB}}$. The conditional entropy $H(f_{\textit{VB}}|f_{\textit{MVS}})$ and $H(f_{\textit{MVS}}|f_{\textit{VB}})$ are calculated as follows:
@@
Figure 6: Visualization results of reconstructed frame using different feature types at $bpp\approx 0.08$. Relying only on pixel-oriented feature $f_{\textit{pixel}}$ tends to miss fine details of the dog’s fur textures, while using only semantic-level feature $f_{\textit{semantic}}$ leads to the loss of grass texture and introduces artifacts. By fusing $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$, our method reconstructs the dog’s fur details better and avoids artifacts in the grass.
@@ III-B3 Discussion
Through the BiEC mechanism, the semantic decoder is encouraged to reconstruct semantic representations by relying only on features that are beneficial for video understanding, guided by the visual backbone. In other words, the intermediate features $f_{\textit{MVS}}$ of the semantic decoder are expected to be semantic-aware. To illustrate this property, we provide a visualization. Specifically, 400 images from 8 classes of the ImageNet-1k [21] validation set are fed into our framework, and the intermediate features $f_{\textit{MVS}}^{1/8}$ of the semantic decoder are projected into a two-dimensional space using t-SNE [51]. Since our framework is designed for P-frames and cannot be evaluated on a single frame, the reference frame is set to all zeros for this visualization. As shown in Fig. 5, the semantic decoder, under the joint effect of the visual backbone and the BiEC mechanism, exhibits improved capability in forming compact and semantically meaningful feature representations.
@@ III-C Semantic-Pixel Dual-Path Fusion
The BiEC-based semantic decoder focuses on features relevant to machine vision tasks, while the motion- and context-related modules aim to reconstruct pixel-level details. These two objectives are not fully aligned. Although machine-oriented tasks and traditional distortion metrics (e.g., PSNR, MS-SSIM) both emphasize structural and color preservation, they differ in their sensitivity to low-level textures and patterns. To bridge this gap, we propose the SPDF module, which injects pixel-level priors into the machine-oriented reconstruction by selectively combining pixel-level features $f_{\textit{pixel}}$ and semantic-level features $f_{\textit{MVS}}$. Specifically, as illustrated in Fig. 6, using only $f_{\textit{MVS}}$ helps retain the texture of the dog’s fur and may increase its detection accuracy, but it ignores the texture of the grass and introduces visible artifacts. In contrast, relying solely on $f_{\textit{pixel}}$ leads to blur under limited bitrates. By gated fusing both $f_{\textit{MVS}}$ and $f_{\textit{pixel}}$, our method effectively suppresses grass artifacts and preserves fine textures of fur.
The structure of the SPDF module is shown in Fig. 3(b). Specifically, the machine-oriented feature $f_{\textit{MVS}}$ and pixel-level features $f_{\textit{pixel}}$ are both fed into the SPDF module. A residual block is employed to calculate an element-wise gating coefficient $\alpha$, which is formulated as follows:
@@
where $U(\cdot)$ denotes U-Net(s), $Conv2d(\cdot)$ denotes a 2D-convolutional layer, $\alpha$ denotes the extracted coefficient map for feature fusion, and $c_{1}$ denotes motion condition [29].
@@ C $D_{\textit{MSE}}(x,\bar{x}),R_{c}$ 1e-4 6 8
Using pixel-domain distortion metrics (e.g., MSE) rather than perception-related ones is essential in the early stage of codec training [4]. Meanwhile, existing optical flow-based inter-prediction methods are typically optimized for pixel-domain quality instead of MVS. Therefore, we divide the training process into two main stages. As shown in Table I, a multi-stage optimization strategy is adopted to achieve optimal rate–task performance.
In the base stage, the network is trained to achieve accurate inter-prediction and fine-grained pixel-domain reconstruction. This stage includes five sub-stages that separately optimize the motion-related and contextual-coding-related modules for improving overall rate-distortion (RD) performance, following [37, 29, 4]. To enhance perceptual quality, learned perceptual image patch similarity (LPIPS) [52] is also introduced, making the baseline codec more suitable for perception-driven tasks. After this stage, the codec is capable of completing the full compression loop. The rate-distortion loss $\mathcal{L}_{\textit{base}}$ of this stage is formulated as follows:
@@ IV-A1 Settings of Vision Models
For the BiEC mechanism, Swin Transformer (base) [53] is employed during the process of training to align information with the proposed semantic decoder. Specifically, the intermediate features from its second, third, and fourth stages are utilized in our framework and named $f_{\textit{VB}}^{1/4}$, $f_{\textit{VB}}^{1/8}$, and $f_{\textit{VB}}^{1/16}$. To maintain machine-perceptual consistency, $res2$ and $res3$ from both ResNet-18 [21] and Swin Transformer (tiny) [53] are employed, along with the $4$-th and $8$-th layers of DINO-v2 (small) [22]. Deeper layers in VFM are excluded to prevent overfitting [8]. Additionally, the reconstructed frames are $2\times$ down-sampled by linear interpolation before being fed into Spynet [35], to prevent overemphasis on pixel-level information during training.
@@ IV-A3 Training Settings
As shown in TABLE I, our training process consists of 50 epochs, with 40 epochs in the base stage and 10 epochs in the VCM stage. The training is conducted on four parallel NVIDIA RTX 3090 GPUs with a global batch size of 8, taking about 20 days. Notably, when training with long GOP, we retain gradients across continuous P-frames to better suppress the cumulative errors introduced by the lossy compression loop. Weight $\lambda_{\textit{MSE}}$ is randomly sampled from $[64,512]$. Since $D_{\textit{LPIPS}}$ and $\mathcal{L}_{\textit{cons}}$ are considered as distortion loss, their weights $\lambda_{\textit{LPIPS}}$ and $\lambda_{\textit{cons}}$ are set to the same value as $\lambda_{\textit{MSE}}$. The weight of the BiEC mechanism $\lambda_{e}$ is set to $1$.
@@ IV-B1 Object Detection
Faster R-CNN [55] is implemented through the Detectron2 codebase [56] and tested on the YTVIS2019 dataset [57], which contains 2,883 videos in 40 categories, ranging in resolution from 360p to 1080p. Since Faster R-CNN has no officially released weights for the YTVIS2019 dataset. We fine-tune pretrained Faster R-CNN on the YTVIS2019 dataset using the default configuration [^1] Configuration of Faster R-CNN-R50: https://github.com/facebookresearch/detectron2/blob/main/configs/COCO-Detection/faster_rcnn_R_50_FPN_1x.yaml for the following experiments.
@@ IV-B4 Multiple Object Tracking
We employ officially released ByteTrack [62] on the MOT17 dataset [63], which contains 42 long videos with frame-level annotations and resolutions ranging from 480p to 1080p.
@@ IV-D Evaluation Metrics
BD-Rate [67] is used to evaluate the rate-task performance of the image and video codec. It measures the average bitrate difference required to achieve the same task performance relative to an anchor codec, where a lower value indicates better compression efficiency. VTM-23.1 serves as the anchor for calculating BD-Rate. In addition, bits per pixel (bpp) reflects the bitrate consumption of videos. A lower bpp value means a higher compression ratio. Task performance is evaluated by the metrics of each downstream task.
@@
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 MOTA IDF IDs [^4] Experiments show that there is a weak correlation between the IDs metric and bitrate, as detailed in Fig. 12. Therefore, the IDs-based BD-rate does not well reflect the rate-task performance. J&F J F
@@ IV-E1 Object Detection
The average precision (AP) under three thresholds [55] is used to evaluate the performance of object detection. The proposed SEC-VCM achieves the best rate-task performance, as shown in TABLE II. It reduces $46.22\%$ (AP), $44.35\%$ (AP50), and $47.55\%$ (AP75) bitrate compared to VTM. The detection performance of all compared methods across various bitrates is shown in Fig. 7. The SEC-VCM consistently outperforms all methods in terms of detection AP, AP50, and AP75 at all tested bitrate points between $bpp=0.067$ and $bpp=0.211$.
@@
Figure 10: Visualization on the YTVIS2019 dataset of video instance segmentation results. Under intense motion (1st row) and instance overlap (2nd row), Mask2Former [59] achieves more accurate instance contours on the reconstructed video from SEC-VCM.
@@ IV-E3 Multiple Object Tracking
Multiple object tracking accuracy (MOTA), ID F1 score (IDF1), and ID switches (IDs) are used to evaluate the performance of tracking, referring to [62]. As shown in TABLE II, our proposed SEC-VCM outperforms other methods in MOTA and achieves $44.9\%$ bitrate saving compared to VTM. Since IDF and IDs are mainly influenced by the inter-frame consistency rather than the frame quality, they show weaker correlation with bitrate. Although SEC-VCM does not have the best rate-task performance in terms of IDF1 and IDs, it outperforms other methods on variable bitrates, as shown in Fig. 12.
@@ IV-E4 Video Object Segmentation
Three metrics—Jaccard value (J), F-measure (F), and J&F—are used to evaluate the performance of video object segmentation. As demonstrated in TABLE II and Fig. 11, our proposed SEC-VCM achieves the most bitrate saving compared to VTM in terms of J&F ($29.83\%$) and F ($38.42\%$) while achieving a comparable compression ratio with SMC++ on the metric of J.
@@ IV-E5 Discussion
The above experiments show that the SEC-VCM achieves promising rate-task performance in most metrics. It is worth mentioning that although many methods exhibit excellent rate-distortion (PSNR, MS-SSIM) and rate-perception (LPIPS, etc.) performance that far exceeds standard codecs, their performance in downstream machine vision tasks is not satisfactory. These widely used metrics cannot well reflect the needs of machine vision tasks [68], which further illustrates the necessity of VCM-related works.
@@ IV-F1 Ablation on the Guidance Visual Backbone in BiEC Mechanism
TABLE IV: Influence of the BiEC loss weight $\lambda_{e}$. $H$ denotes the value of conditional entropy. VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$. The best rate-task performance is displayed in red.
@@
$M_{7}$ KL KL +17.3%
@@ IV-F2 Ablation on the BiEC Mechanism
We conduct ablation studies on the bi-directional design and the multi-layer design, as shown in Table V. Results from $M_{1}$ and $M_{2}$ demonstrate that each direction of the entropy constraints plays a critical role in improving rate-task performance (average 11.7%). Also, the experimental results of $M_{3}$, $M_{4}$, and $M_{5}$ show that removing multi-layer design results in significant rate-task performance degradation (average 12.0%). Notably, in $M_{6}$, the proposed conditional entropy loss $\mathcal{L}_{e}$ is replaced with MSE, leading to a 21.1% decrease in rate-task performance, highlighting the superiority of soft alignment over rigid matching.
@@
PLVC (IJCAI’22) [40] [^7] Since the officially released code of PLVC is not supported by the NVIDIA RTX 3090 and the corresponding CUDA version, it is tested on CPU. The discriminator from its GAN architecture is not included in the number of parameters. / / / 19.2M
@@
Figure 13: Visualization of the intermediate results in the gated fusion process. The pixel-oriented feature $f_{\textit{pixel}}$ and machine-oriented feature $f_{\textit{semantic}}$ are selectively combined based on $\alpha$ for the final reconstruction. Red and blue indicate high and low feature values, respectively.

### v2

@@ III-B2 Bi-directional Entropy-Constraint Mechanism
Figure 5: The t-SNE visualization of intermediate features $f_{\textit{MVS}}^{1/8}$ in the semantic decoder without and with symmetric information alignment. After alignment via the BiEC mechanism, intra-class distances are effectively reduced, and the intermediate features $f_{\textit{MVS}}^{1/8}$ in the decoding process exhibit more discriminative semantic representations.
To enforce symmetry, we further impose a conditional entropy constraint in the reverse direction, thus making $f_{\textit{VB}}$ a subset of $f_{\textit{MVS}}$. Through this bi-directional constraint, the features $f_{\textit{MVS}}$ in the semantic decoder and the visual-backbone-extracted representations $f_{\textit{VB}}$ are softly aligned, as illustrated in Fig. 4. The above process is formulated as follows:
@@
Specifically, each DE module contains a convolutional layer and a ResBlock [24]. Estimated distribution parameters $\mu$ and $\sigma$ have the same shape as the corresponding feature $f$. Following existing methods [46, 31], the Laplace distribution is used to model the conditional probability of the features $f_{\textit{MVS}}$ and $f_{\textit{VB}}$. The conditional entropy $H(f_{\textit{VB}}|f_{\textit{MVS}})$ and $H(f_{\textit{MVS}}|f_{\textit{VB}})$ are calculated as follows:
@@
Figure 6: Visualization results of reconstructed frame using different feature types at $bpp\approx 0.08$. Relying only on pixel-oriented feature $f_{\textit{pixel}}$ tends to miss fine details of the dog’s fur texture, while using only semantic-level feature $f_{\textit{semantic}}$ leads to the loss of grass texture and introduces artifacts. By fusing $f_{\textit{semantic}}$ and $f_{\textit{pixel}}$, our method reconstructs the dog’s fur details better and avoids artifacts in the grass.
@@ III-B3 Discussion
Through the BiEC mechanism, the semantic decoder is encouraged to reconstruct semantic representations by relying only on features that are beneficial to video understanding, guided by the visual backbone. In other words, the intermediate features $f_{\textit{MVS}}$ of the semantic decoder are expected to be semantic-aware. To illustrate this property, we provide a visualization. Specifically, 400 images from 8 classes of the ImageNet-1k [24] validation set are fed into our framework, and the intermediate features $f_{\textit{MVS}}^{1/8}$ of the semantic decoder are projected into a two-dimensional space using t-SNE [47]. Since our framework is designed for P-frames and cannot be directly evaluated on single frames, all pixels of the reference frame are set to zero for this visualization. As shown in Fig. 5, the semantic decoder, under the joint effect of the visual backbone and the BiEC mechanism, exhibits improved capability in forming compact and semantically meaningful feature representations.
@@ III-C Semantic-Pixel Dual-Path Fusion
The BiEC-based semantic decoder focuses on features relevant to machine vision tasks, while the motion- and context-related modules aim to reconstruct pixel-level details. These two objectives are not fully aligned. Although machine-oriented tasks and traditional distortion metrics (e.g., PSNR, MS-SSIM) both emphasize structural and color preservation, they differ in their sensitivity to low-level textures and patterns. To bridge this gap, we propose the SPDF module, which injects pixel-level priors into the machine-oriented reconstruction by selectively combining pixel-level features $f_{\textit{pixel}}$ and semantic-level features $f_{\textit{MVS}}$. Specifically, as illustrated in Fig. 6, using only $f_{\textit{MVS}}$ helps retain the texture of the dog’s fur and may increase its detection accuracy, but it ignores grass textures and introduces artifacts, which undermine the semantics of the grass region. In contrast, relying solely on $f_{\textit{pixel}}$ leads to a blur under limited bitrates. By fusing both $f_{\textit{MVS}}$ and $f_{\textit{pixel}}$, our method effectively suppresses grass artifacts and preserves fine textures of fur.
The structure of the SPDF module is shown in Fig. 3(b). Specifically, the machine-oriented feature $f_{\textit{MVS}}$ and pixel-level feature $f_{\textit{pixel}}$ are both fed into the SPDF module. A residual block is employed to compute an element-wise gating coefficient $\alpha$, which is formulated as follows:
@@
where $U(\cdot)$ denotes U-Net(s), $Conv2d(\cdot)$ denotes a 2D-convolutional layer, $\alpha$ denotes the extracted coefficient map for feature fusion, and $c_{1}$ denotes the motion condition [31].
@@ C $D_{\textit{MSE}}(x,\bar{x}),R_{c}$ 1e-4 6 8
Using pixel-domain distortion metrics (e.g., MSE) rather than perception-related metrics is essential in the early stage of codec training [4]. Meanwhile, existing optical flow-based inter-prediction methods are typically optimized for pixel-domain quality instead of MVS. Therefore, we divide the training process into two main stages. As shown in Table I, a multi-stage optimization strategy is adopted to achieve optimal rate-task performance.
In the base stage, the network is trained to achieve accurate inter-prediction and fine-grained pixel-domain reconstruction. This stage includes five sub-stages that separately optimize the motion-related and contextual-coding-related modules to improve overall rate-distortion (RD) performance, following [36, 31, 4]. To enhance perceptual quality, learned perceptual image patch similarity (LPIPS) [48] is also introduced, making the baseline codec more suitable for perception-driven tasks. After this stage, the codec is capable of completing the full compression loop. The rate-distortion loss $\mathcal{L}_{\textit{base}}$ of this stage is formulated as follows:
@@ IV-A1 Settings of Vision Models
For the BiEC mechanism, Swin Transformer (base) [49] is employed during the process of training to align information with the proposed semantic decoder. Specifically, the intermediate features from its second, third, and fourth stages are utilized in our framework and named $f_{\textit{VB}}^{1/4}$, $f_{\textit{VB}}^{1/8}$, and $f_{\textit{VB}}^{1/16}$. To maintain machine-perceptual consistency, $res2$ and $res3$ from both ResNet-18 [24] and Swin Transformer (tiny) [49] are employed, along with the $4$-th and $8$-th layers of DINO-v2 (small) [25]. Deeper layers in VFM are excluded to prevent overfitting [9]. Additionally, the reconstructed frames are twice down-sampled by linear interpolation before being fed into Spynet [34], to prevent overemphasis on pixel-level information during training.
@@ IV-A3 Training Settings
As shown in TABLE I, our training process consists of 50 epochs, with 40 epochs in the base stage and 10 epochs in the VCM stage. The training is conducted on four parallel NVIDIA RTX 3090 GPUs with a global batch size of 8, taking about 20 days. Notably, when training with long GOP, we retain gradients across continuous P-frames to better suppress the cumulative errors introduced by the lossy compression loop. Weight $\lambda_{\textit{MSE}}$ is randomly sampled from $[64,512]$. Since $D_{\textit{LPIPS}}$ and $\mathcal{L}_{\textit{cons}}$ are considered distortion losses, their weights $\lambda_{\textit{LPIPS}}$ and $\lambda_{\textit{cons}}$ are set to the same value as $\lambda_{\textit{MSE}}$. The weight of the BiEC mechanism $\lambda_{e}$ is set to $1$.
@@ IV-B1 Object Detection
Faster R-CNN [51] is implemented through the Detectron2 codebase [52] and tested on the YTVIS2019 dataset [53], which contains 2,883 videos in 40 categories, ranging in resolution from 360p to 1080p. Since Faster R-CNN has no officially released weights on the YTVIS2019 dataset. We fine-tune pretrained Faster R-CNN on the YTVIS2019 dataset using the default configuration [^1] Configuration of Faster R-CNN-R50: https://github.com/facebookresearch/detectron2/blob/main/configs/COCO-Detection/faster_rcnn_R_50_FPN_1x.yaml for the following experiments.
@@ IV-B4 Multiple Object Tracking
We employ officially released ByteTrack on the MOT17 dataset [58], which contains 42 long videos with frame-level annotations and resolutions ranging from 480p to 1080p.
@@ IV-D Evaluation Metrics
BD-Rate [63] is used to evaluate the rate-task performance of the image and video codecs. It measures the average bitrate difference required to achieve the same task performance relative to an anchor codec, where a lower value indicates better compression efficiency. VTM-23.1 serves as the anchor for calculating BD-Rate. In addition, bits per pixel (bpp) reflects the bitrate consumption of videos. A lower bpp value means a higher compression ratio. Task performance is evaluated by the metrics of each downstream task.
@@
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75 MOTA IDF1 IDs [^4] Experiments show that there is a weak correlation between the IDs metric and bitrate, as detailed in Fig. 12. Therefore, the IDs-based BD-rate does not well reflect the rate-task performance. J&F J F
@@ IV-E1 Object Detection
The average precision (AP) under three thresholds [51] is used to evaluate the performance of object detection. The proposed SEC-VCM achieves the best rate-task performance, as shown in TABLE II. It reduces $46.2\%$ (AP), $44.4\%$ (AP50), and $47.6\%$ (AP75) bitrate compared to VTM. The detection performance of all compared methods across various bitrates is shown in Fig. 7. The SEC-VCM consistently outperforms all methods in terms of detection AP, AP50, and AP75 at all tested bitrate points between $bpp=0.067$ and $bpp=0.211$.
@@
Figure 10: Visualization on the YTVIS2019 dataset of video instance segmentation results. Under intense motion (1st row) and instance overlap (2nd row), Mask2Former [55] achieves more accurate instance contours on the reconstructed videos from SEC-VCM.
@@ IV-E3 Multiple Object Tracking
Multiple object tracking accuracy (MOTA), ID F1 score (IDF1), and ID switches (IDs) are used to evaluate the performance of tracking, referring to [58]. As shown in TABLE II, our proposed SEC-VCM outperforms other methods in MOTA and achieves $44.9\%$ bitrate saving compared to VTM. Since IDF1 and IDs are mainly influenced by the inter-frame consistency rather than the frame quality, they show weaker correlation with bitrate. Although SEC-VCM does not have the best rate-task performance in terms of IDF1 and IDs, it outperforms other methods on variable bitrates, as shown in Fig. 12.
@@ IV-E4 Video Object Segmentation
Three metrics—Jaccard value (J), F-measure (F), and J&F—are used to evaluate the performance of video object segmentation. As demonstrated in TABLE II and Fig. 11, our proposed SEC-VCM achieves the most bitrate saving compared to VTM in terms of J&F ($29.8\%$) and F ($38.4\%$) while achieving a comparable compression ratio with SMC++ on the J metric.
@@ IV-E5 Discussion
The above experiments show that the SEC-VCM achieves promising rate-task performance in most metrics. It is worth mentioning that although many methods exhibit excellent rate-distortion (PSNR, MS-SSIM) and rate-perception (LPIPS, etc.) performance that far exceeds standard codecs, their performance in downstream machine vision tasks is not satisfactory. These widely used distortion metrics cannot well reflect the needs of machine vision tasks, which further illustrates the necessity of VCM-related works. Also, developing general and differentiable metrics tailored for downstream tasks remains an urgent and open problem [64].
@@ IV-F1 Ablation on the Guidance Visual Backbone in BiEC Mechanism
TABLE IV: Influence of the BiEC loss weight $\lambda_{e}$. $H$ denotes the value of conditional entropy. VTM-23.1 is the anchor for calculating BD-rate (%) $\downarrow$. The best rate-task performance is in red.
@@
$M_{7}$ $\text{KL}_{\text{channel}}$ $\text{KL}_{\text{channel}}$ +17.3%
$M_{8}$ $\text{KL}_{\text{spatial}}$ $\text{KL}_{\text{spatial}}$ +13.5%
@@ IV-F2 Ablation on the BiEC Mechanism
We conduct ablation studies on the bi-directional design and the multi-layer design, as shown in Table V. Results from $M_{1}$ and $M_{2}$ demonstrate that each direction of the entropy constraints plays a critical role in improving rate-task performance of 11.7%. Also, the experimental results of $M_{3}$, $M_{4}$, and $M_{5}$ show that removing multi-layer design results in significant rate-task performance degradation of 12.0%.
Notably, in $M_{6}$, the proposed conditional entropy loss $\mathcal{L}_{e}$ is replaced by MSE to perform direct element-wise alignment, resulting in a 21.1% decrease in rate-task performance. We also applied channel-wise ($M_{7}$) and spatial-wise ($M_{8}$) KL divergence losses to achieve distribution-level alignment, which led to performance drops of 17.3% and 13.5%, respectively. These results highlight the superiority of the proposed information-level alignment over direct (MSE) and distribution-level (KL divergence) alignment.
@@
PLVC (IJCAI’22) [37] [^7] Since the officially released code of PLVC is not supported by the NVIDIA RTX 3090 and the corresponding CUDA version, we tested it on CPU. Also, the discriminator in its GAN architecture is not included in the number of parameters. / / / 19.2M
@@
Figure 13: Visualization results of the gated fusion process. Red and blue indicate high and low feature values, respectively. The pixel-oriented feature $f_{\textit{pixel}}$ and machine-oriented feature $f_{\textit{semantic}}$ are selectively combined based on $\alpha$ for the final reconstruction.